OpenStudy (anonymous):

Help prove that (cosx+cotx)/(1+cscx)=cosx ??

zepdrix (zepdrix):

\[\Large\bf\sf \frac{\cos x+\cot x}{1+\csc x}\quad=\quad \cos x\]Can you see the fancy math up above^? Or is it just a big jumbled mess? LaTeX plugin hasn't been working well today, just checking first.

OpenStudy (anonymous):

It looks kinda funky...haha

zepdrix (zepdrix):

grr :C

zepdrix (zepdrix):

Ahhh I think I finally got it. This was a tricky problem :U

zepdrix (zepdrix):

The draw tool won't work either.. lemme uhhh.. upload a picture to show you.. one sec.

OpenStudy (anonymous):

wow that cleared it up a lot! Thanks!